OpenStudy (anonymous):

Solve for x: 2 over 3 (x - 4) = 2x. 2 -2 -8 -4

HanAkoSolo (jamierox4ev3r):

\(\large\frac{2}{3}\)(x-4)=2x just multiply everything in the parentheses by 2/3 , which will give you 2/3x -8/3

HanAkoSolo (jamierox4ev3r):

@MyChem then you would solve for x like a regular equation. THen you would get x=-2. to check this answer, plug in the value to replace x and then you shall have your answer

HanAkoSolo (jamierox4ev3r):

to verify whether it is correct or not

OpenStudy (anonymous):

would it be -2?

OpenStudy (anonymous):

Solve: 4(x + 1) = 4x + 10 0 6 All Real Numbers No Solution

HanAkoSolo (jamierox4ev3r):

well just by looking at this, it looks like this is saying 4x+4=4x+10, which is not true. Therefore, this is NO SOLUTION

OpenStudy (anonymous):

if I give you a medal will you help me with mah other questions? ^^

HanAkoSolo (jamierox4ev3r):

of course :)

HanAkoSolo (jamierox4ev3r):

well if L=24 W=11 units, then 24+24+11+11=70. So no it cannot because x=48+2y

OpenStudy (anonymous):

which do you think it would be? o.o

OpenStudy (anonymous):

nvm bout that question The table below shows two equations: Equation 1 |2x – 3| + 5 = 4 Equation 2 |5x + 3| − 10 = 3 Which statement is true about the solution to the two equations? Equation 1 and equation 2 have no solutions. Equation 1 has no solution and equation 2 has solutions x = 2, -3.2. The solutions to equation 1 are x = 1, 2 and equation 2 has no solution. The solutions to equation 1 are x = 1, 2 and equation 2 has solutions x = 2, -3.2.